OpenStudy (anonymous):

Which function represents a translation of the graph of y=x^2 by 8 units to the right? A. y=(x+8)^2 B. y=x^2+8 C. y=(x-8)^2 D. y=8x^2

OpenStudy (anonymous):

Do you have an idea already? Or are you not sure how to figure it out?

OpenStudy (anonymous):

i think its b but i wanted some outside confirmation feel me?

OpenStudy (anonymous):

Sure. It's not b.

OpenStudy (anonymous):

welll then i have no idea

OpenStudy (anonymous):

In general, lets say you are looking at translations of a function: y = x^2 The translations will be of the form: y-k = (x-h)^2 You can add k to both sides so it might look more like: y = (x-h)^2 + k But in anycase, the h represents a shift in the left/right direction, the k is a shift up and down.

OpenStudy (anonymous):

So with that in mind, do you have another guess?

OpenStudy (anonymous):

oh i see so the answer should be c?

OpenStudy (anonymous):

oh i see so the answer should be c?

OpenStudy (anonymous):

oh i see so the answer should be c?

OpenStudy (anonymous):

oh i see so the answer should be c?

OpenStudy (lalaly):

f(x) = x^2 f(x) = x^2 + 8 <---that shifts the function up 8 units on the graph f(x) = x^2 - 8 <----that shifts the function down 8 units on the graph Shifting right or left is a little more tricky, because you have to directly modify the function's variable. For example: f(x) = x^2 f(x+8) = (x+8)^2 <----that will shift the graph 8 points to the left (its inverse for horizontal translation, positive = to the left negative = to the right) f(x-8) = (x-8)^2 <----that will shift the graph 8 points to the right

OpenStudy (anonymous):

oh i see so the answer should be c?

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

If it were y=(x+8)^2 it would be a shift of 8 to the left instead of the right because it's y=(x-(-8))^2